OpenStudy (anonymous):

Solve for d. (Question in answers) A. 4, –4 B. 5, –5 C. 5, 6 D. 6, –6

OpenStudy (anonymous):

\[\sqrt{D ^{2} -11} = 5\]

OpenStudy (anonymous):

@Mertsj :)

OpenStudy (mertsj):

Square both sides.

OpenStudy (mertsj):

What do you get?

OpenStudy (anonymous):

\[\sqrt{D - 11}^{2} = 5^{?}\]

OpenStudy (mertsj):

\[(\sqrt{d^2-11})^2=5^2\]

OpenStudy (mertsj):

Can you simplify that?

OpenStudy (anonymous):

becuase we multi ply d by d, then find the sqrrt isnt it just D?

OpenStudy (anonymous):

5x5 25

OpenStudy (mertsj):